OpenStudy (anonymous):

How do I simplify the expression 4+√-4+i i being an imaginary number.

OpenStudy (anonymous):

Is that \[4 + \sqrt{-4} + i\] or \[4 + \sqrt{-4 + i}\]

OpenStudy (anonymous):

The first one

OpenStudy (anonymous):

Oh ok, thank you. Would you agree that \[\sqrt{-4} = \sqrt{(-1)(4)} = \sqrt{-1} * \sqrt{4}\]

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

And \[\sqrt{-1} = i\]

OpenStudy (anonymous):

Yes, isnt also i^2 -1

OpenStudy (anonymous):

That is true as well

OpenStudy (anonymous):

So we see how \[\sqrt{-4} = \sqrt{-1} * \sqrt{4} = i * 2 = 2i\]

OpenStudy (anonymous):

Ohh I see

OpenStudy (anonymous):

So would my answer be 4+3i ?

OpenStudy (anonymous):

Yes, that is right

OpenStudy (anonymous):

Oh Ok, Thank you!